There is a bag with 7 blue marbles and 4 orange marbles. You pull a marble out of the bag three times, without replacing the marble each time. What is the probability that you will pull out three blue marbles?
Answer #1

probability that you will pull out three blue marbles =P(select 3 blue from 7) =(7C3)(11C3) =(7/11)*(6/10)*(5/9)=7/33